### v1

# Meta Pruning via Graph Metanetworks : A Meta Learning Framework for Network Pruning

Abstract
Network pruning, aimed at reducing network size while preserving accuracy, has attracted significant research interest. Numerous pruning techniques have been proposed over time. They are becoming increasingly effective, but more complex and harder to interpret as well. Given the inherent complexity of neural networks, we argue that manually designing pruning criteria has reached a bottleneck. To address this, we propose a novel approach in which we use a neural network to prune neural networks. More specifically, we introduce the newly developed idea of metanetwork from meta-learning into pruning. A metanetwork is a network that takes another network as input and produces a modified network as output. In this paper, we first establish a bijective mapping between neural networks and graphs, and then employ a graph neural network as our metanetwork. We train a metanetwork that learns the pruning strategy automatically which can transform a network that is hard to prune into another network that is much easier to prune. Once the metanetwork is trained, our pruning needs nothing more than a feedforward through the metanetwork and the standard finetuning to prune at state-of-the-art. Our method achieved outstanding results on many popular and representative pruning tasks (including ResNet56 on CIFAR10, VGG19 on CIFAR100, ResNet50 on ImageNet). Our code is available at https://github.com/Yewei-Liu/MetaPruning.
1 Introduction
Correspondence to Muhan Zhang
With the rapid advancement of deep learning [33, 53], neural networks have become increasingly powerful. However, this improved performance often comes with a significant increase in the number of parameters and computational cost (FLOPs). As a result, there is growing interest in methods to simplify these networks while preserving their performance. Pruning, which involves selectively removing certain parts of a neural network, has proven to be an effective approach [2, 17, 51].
A key idea of a large number of previous pruning works is to remove the unimportant components of a neural network. To achieve this goal, various criteria that measure the importance of components of a neural network have been invented. Some criteria are based on norm or magnitude scores [15, 36, 55, 18]. Some are based on how much the performance changes when certain parts of the network are removed or preserved [68, 46, 10, 66]. Some others defined more complex criteria such as saliency and sensitivity [32, 34, 72].
Once a pruning criterion is established, we can assess whether a model is easy or difficult to prune based on that criterion. For instance, using a simple $l_{2}$ norm criterion, a model whose parameters exhibit either very large or very small $l_{2}$ norms is clearly easier to prune compared to one with uniformly distributed $l_{2}$ norms. From this viewpoint, we can classify models as more or less suitable for pruning. This concept forms the foundation of our work.
Previously, researchers have generally approached pruning improvement in two ways: either by refining the pruning criteria or by transforming the model from “bad” to “good” according to a given criterion. The first approach is more straightforward and has been extensively studied in prior work. In contrast, the second approach has received comparatively less attention. However, several studies adopting this latter perspective have achieved promising results. Notably, sparsity regularization methods [61, 9, 60] exemplify this approach by encouraging models to become sparser, thereby facilitating more effective pruning.
Our work follows the second approach and introduces a novel method inspired by sparsity regularization. We focus on this approach because, despite the many pruning criteria proposed, most lack solid theoretical foundations and have become overly complex. Therefore, in this work, we adopt a simple variant of the $l_{2}$ norm—the group norm from Fang et al. [9]—as our pruning criterion. Additionally, we bring the concept of metanetworks from meta-learning into pruning. A metanetwork is a network that takes another neural network as input and produces a modefied neural network as output. In our method, the metanetwork transforms a “bad” network into a “good” one, making it more amenable to pruning. Once trained, our metanetwork-based approach requires only a feedforward pass through the metanetwork and some standard finetuning to achieve state-of-the-art pruning results.
Our contributions can be summarized as follows:
  • Introducing the idea of metanetwork in meta-learning into pruning for the first time.
  • Proposing a novel pruning method that is simple and powerful.
  • Achieving state-of-the-art performance on several popular pruning tasks.
It is worth noting that this paper focuses solely on structural pruning, which does not rely on specialized AI accelerators or software to reduce memory usage and computational costs. This makes it more versatile and broadly applicable. However, all our methods can also be applied directly to unstructured pruning.
2 Related Work
2.1 Metanetworks
Using neural networks to process other neural networks has emerged as an intriguing research direction. Early studies demonstrated that neural networks can extract useful information directly from the weights of other networks [58]. We define a metanetwork as a neural network that takes another neural network as input and outputs either information about it or a modified network. Initial metanetworks were simple multilayer perceptrons (MLPs), while more recent designs incorporate stronger inductive biases by preserving symmetries [13] inherent in the input networks. Broadly, metanetwork architectures can be categorized into two perspectives. The weight space view applies specially designed MLPs directly on the model’s weights [71, 48, 73, 57, 74, 75, 56]. In contrast, the graph view transforms the input network into a graph and applies graph neural networks (GNNs) to it [40, 30, 27]. In this work, we adopt the graph view to build our metanetwork, leveraging the natural correspondence between graph nodes and neurons in a network layer, as well as the symmetries they exhibit [47]. Given the maturity and effectiveness of GNNs, the graph metanetwork approach offers both elegance and practicality.
2.2 Graph Neural Networks
Graph neural networks (GNNs) are a class of neural networks specifically designed to operate on graph-structured data, where graphs consist of nodes and edges with associated features [63, 52, 29]. GNNs have gained significant attention in recent years, with frameworks such as the message passing neural network (MPNN) [12] proving to be highly effective. In this paper, we employ the message passing framework, specifically using Principal Neighborhood Aggregation (PNA) [3] as the backbone architecture for our metanetwork.
2.3 Sparsity Regularization-Based Pruning
Networks that are inherently sparse tend to be easier to prune effectively. Prior research has explored various regularization techniques to encourage sparsity in neural networks to facilitate pruning. This idea of sparsity regularization has been widely adopted across many works [61, 14, 25, 44, 22, 43, 64, 65, 9, 60]. Our method inherits this principle to some extent, as our metanetwork can be viewed as transforming the original network into a sparser version that is more amenable to pruning.
2.4 Learning to Prune & Meta Pruning
Given the complexity of pruning, several previous works have leveraged neural networks [5, 22, 62, 1], reinforcement learning [50, 70, 23], and other techniques to automatically learn pruning strategies. Viewing pruning as a learning problem naturally leads to the idea of meta-learning, which learns “how to learn” [24]. Some prior works have specifically applied meta-learning to pruning [45, 37]. While our work also adopts a meta-learning framework, it differs fundamentally from these previous approaches by using a metanetwork.
3 Method
Our approach centers on training a metanetwork that learns the pruning strategy automatically and can transform a neural network that is difficult to prune into one that is easier to prune, thereby improving pruning effectiveness and efficiency.
We begin by describing how our metanetwork transforms one network into another, including the conversion between network and graph, and the design of our metanetwork(a graph neural network). Next, we introduce the group $l_{2}$ norm criterion for meta-training and pruning. Then we detail the meta-training pipeline used to train the metanetwork. Finally, we present the complete pruning process.
[…]
3.1 Equivalent Conversion between Neural Network and Graph
To enable a neural network to be processed by another network, we first need to represent the original network in a structured data format. In this work, we convert the neural network into a graph, drawing inspiration from prior works such as Kofinas et al. [30], Lim et al. [40], but with notable differences in implementation.
To convert a network into a graph, we establish the correspondence between the components of a neural network and elements of a graph as follows:
  • Node: Each neuron or convolutional channel in the network is represented as a node in the graph.
  • Edge: An edge exists between two nodes if there is a direct connection between them in the original network. This includes connections between neurons or channels within linear or convolutional layers, as well as residual connections.
  • Node Features: Node features comprise parameters associated with neurons, here we use the weight, bias, running mean, running var of batchnorm [26], including both batchnorms from previous adjacent layers and previous skip connection layers (if exist).
  • Edge Features: Edge features encode the weights of connections between nodes. We treat linear connections and residual connections as special cases of convolutional connections with a kernel size of 1. For a $k\times k$ convolution between two channels, we linearize it into a $k^{2}$-dim edge feature vector.
Since graph neural networks require all edge features to have the same dimension, we adopt one of two strategies to standardize them:
  • Pad all convolutional kernels to the same size (e.g., the maximum kernel size in the network), then flatten them to form fixed-length edge feature vectors.
  • Flatten convolutional kernels of varying sizes, then apply learned linear transformations to project them into a uniform feature dimension.
By following these rules, we can generate a graph that corresponds to a given neural network. Conversely, the inverse transformation can be applied to convert a graph back into an equivalent network. In essence, we establish an equivalent conversion between networks and graphs. For a more intuitive understanding, see Figure 1.
Figure 1: We visualize a two layer small CNN as example. Our graph contains 6 nodes corresponding with 6 neurons in the network. For neurons connected via convolutional channel like (1,3) and connected by residual connection like (1,5), we add an edge between their corresponding nodes. Then we generate node and edge features as shown in the figure. For some special cases, we use default values to replace non-existent values. For example, for node 3 who doesn’t have a residual connection, we set the last 4 number of its node feature as [1, 0, 0, 1]. This means we treat it as with previous residual batchnorm of weight 1, bias 0, running mean 0, running var 1, performing the same as with no batchnorm.
3.2 Graph metanetwork design
After changing the input network into a graph, we use our metanetwork, a graph neural network, to process the input graph. We build our metanetwork based on the message passing framework [12] and PNA [3] architecture.
We use $v_{i}$ to represent node $i$’s feature vector, and $e_{ij}$ to represent feature vector of the edge between $i$ and $j$. We consider undirected edges, i.e., $e_{ij}$ is the same as $e_{ji}$. Our metanetwork is as follows.
First, all input node and edge features are encoded into the same hidden dimension.
$v\leftarrow\text{MLP}_{NodeEnc}(v^{in}),$ (1)
$e\leftarrow\text{MLP}_{EdgeEnc}(e^{in}).$ (2)
Then they will pass through several message passing layers. In each layer, we generate the messages:
$m_{ij}$ $\leftarrow\text{MLP}_{Node}^{1}(v_{i})\odot\text{MLP}_{Node}^{2}(v_{j})\odot e_{ij},$ (3)
$m^{\prime}_{ij}$ $\leftarrow\text{MLP}_{Node}^{1}(v_{j})\odot\text{MLP}_{Node}^{2}(v_{i})\odot(e_{ij}\odot EdgeInvertor),$ (4)
where $m_{ij}$ and $m^{\prime}_{ij}$ respectively encode message from $i$ to $j$ and message from $j$ to $i$, and EdgeInvertor is defined as:
$EdgeInvertor\triangleq[\underbrace{1,1,\dots,1}_{\text{hidden\_dim}/2},\underbrace{-1,-1,\dots,-1}_{\text{hidden\_dim}/2}].$ (5)
The intuition behind EdgeInvertor is to let the first half dimensions to learn undirectional features invariant to exchanging $i,j$, and the second half to capture directional information that changes equivariantly with reverting $i,j$. Empirically we found this design the most effective among various ways to encode edge features.
Then, we aggregate the messages to update node features:
$v_{i}\leftarrow v_{i}+\text{PNA}_{Aggr}(m_{ij})+\text{PNA}_{Aggr}(m_{ij}^{\prime}),$ (6)
where we have
$\text{PNA}_{Aggr}(m_{ij})\triangleq\text{MLP}_{Aggr}\left(\left[\mathop{\text{MEAN}}\limits_{j:(i,j)\in E}(m_{ij}),\mathop{\text{STD}}\limits_{j:(i,j)\in E}(m_{ij}),\mathop{\text{MAX}}\limits_{j:(i,j)\in E}(m_{ij}),\mathop{\text{MIN}}\limits_{j:(i,j)\in E}(m_{ij})\right]\right).$ (7)
For edge features in each layer, we also update them by:
$e_{ij}\leftarrow e_{ij}$ $+\text{MLP}_{Edge}^{1}(v_{i})\odot\text{MLP}_{Edge}^{2}(v_{j})\odot e_{ij}$ (8)
$+\text{MLP}_{Edge}^{1}(v_{j})\odot\text{MLP}_{Edge}^{2}(v_{i})\odot(e_{ij}\odot EdgeInvertor).$
Finally, we use a decoder to recover node and edge features to their original dimensions:
$v_{pred}\leftarrow\text{MLP}_{NodeDec}(v),$ (9)
$e_{pred}\leftarrow\text{MLP}_{EdgeDec}(e).$ (10)
We multiply our predictions by a residual coefficient and add them to the original inputs as our outputs:
$v_{out}=\alpha\cdot v_{pred}+v_{in},$ (11)
$e_{out}=\beta\cdot e_{pred}+e_{in},$ (12)
where $\alpha$ and $\beta$ in practice are set to a small real numbers like $0.01$. This design enables the metanetwork to learn only the delta weights on top of the original weights, instead of a complete reweighting.
3.3 Group $l_{2}$ Norm Criterion
We use group $l_{2}$ norm from Fang et al. [9] as our pruning criterion. The design philosophy of group $l_{2}$ norm is to measure the sparsity of pruning groups. In structural pruning, we only prune a group of elements in a network. For example, if we prune a neuron, then all edges connected to the neuron should also be pruned, which we say that they are in a group. A more detailed explanation of group can be found in Fang et al. [9]. Since structural pruning is performed on groups, traditional $l_{2}$ norm that treats nodes and edges in isolation is no longer viable. Group $l_{2}$ norm that takes a whole group into consideration has shown to be more effective.
Given a parameter group $g=\{w_{1},w_{2},\ldots,w_{|g|}\}$. For each parameters $w$ with $K$ prunable dimensions indexed by $w[k]$, we measure the importance of the k-th prunable dimension by :
$I_{g,k}=\sum_{w\in g}\|w[k]\|^{2}$ (13)
During training, our sparsity loss is defined as :
$\mathcal{R}(g,k)=\sum_{k=1}^{K}\gamma_{k}\cdot I_{g,k}=\sum_{k=1}^{K}\sum_{w\in g}\gamma_{k}\|w[k]\|^{2}$ (14)
Where $\gamma_{k}$ refers to a shrinkage strength applied to the parameters to modify the gradients for better training. Defined as :
$\gamma_{k}=2^{\alpha(I_{g}^{\max}-I_{g,k})(I_{g}^{\max}-I_{g}^{\min})}$ (15)
And we use $\alpha=4$ in all our experiments.
When pruning, we calculate the normalized relative score of all prunable dimensions in all groups
$\hat{I}_{g,k}=N\cdot I_{g,k}/\sum\{\text{TopN}(I_{g})\}$ (16)
where $N$ is a hyperparameter that chooses the top $N$ importance scores in the group $g$ to normalize every prunable dimension $k$’s score $I_{g,k}$. We prune unimportant dimensions of groups globally from lower to higher according to the normalized relative score.
Figure 2: Meta-Training Pipeline.
3.4 Meta-Training Pipeline
The goal of training our metanetwork to enable it to transform a hard-to-prune network into an easy-to-prune one, referred to as meta-training. This section details our meta-training pipeline, which is illustrated in Figure 2.
To prepare for meta-training, first we need the training data. Meta-training data is comprised of two parts. The first is a set of neural networks to prune, typically between 2 to 8, which serve as the origin network in meta-training. The second is the traditional training datasets, such as CIFAR-10, CIFAR-100, or ImageNet, used to calculate the accuracy loss.
During each training iteration, we select one model from data models as the origin network. The origin network is then converted into its graph representation. This graph is passed through the metanetwork, which outputs a new graph. We subsequently convert this new graph back into a new network.
Two types of losses are computed on the new network:
1) Accuracy Loss: We feed the training data into the new network and compute the cross-entropy loss based on the output predictions, ensuring the new network to have high accuracy after finetuning.
2) Sparsity Loss: We apply the group $l_{2}$ norm (as defined in 14) on the new network to encourage structured sparsity, making it more amenable to pruning.
The gradients from the loss are backpropagated through the metanetwork to update its parameters.
The accuracy loss ensures that the metanetwork preserves the performance of the original network, while the sparsity loss encourages the metanetwork to generate a network with group sparsity that facilitates pruning. After meta-training, our metanetwork can generate a transformed network that has almost no loss in accuracy after finetuning, and is significantly easier to prune.
Empirically, we observe that the metanetwork performs much better on hard-to-prune origin networks. However, most networks are initially easy to prune—for example, ResNet-56 on CIFAR-10 can be pruned with a 1.3× speedup and finetuned without accuracy loss. Therefore, in all our experiments (both during meta-training and pruning), we first prune the network slightly—without any accuracy degradation—to make it harder to prune. This pre-pruning step allows us to fully leverage the capability of the metanetwork.
3.5 Pruning a Target Network
Our pruning pipeline only involves a feedforward through metanetwork and finetuning. It can be summarized as:
$InitialPruning(Finetuning)\rightarrow Metanetwork(Finetuning)\rightarrow Pruning(Finetuning)$
As discussed in 3.4, we first do initial pruning to make our network harder to prune while maintaining the accuracy. Then we feed the hard-to-prune network through the metanetwork to transform it into a much easier-to-prune one, which also has almost no accuracy loss after finetuning. In the end, we perform the final pruning to reduce its size, and the final finetuning to evaluate the performance.
4 Experiments
4.1 Setup
We carry out our experiments on three most classical, popular, and representative image recognition tasks, including resent56 [16] on CIFAR10 [31], VGG19 [54] on CIFAR100 [31] and resnet50 [16] on ImageNet [4]. See appendix A for more general setups.
[…]
4.2 ResNet56 on CIFAR10
Table 1: ResNet56 on CIFAR10 (see full table in appendix 4)
Method Base Pruned $\Delta$ Acc Pruned FLOPs Speed Up
FPGM [19] 93.59 92.93 -0.66 52.6% 2.11
FPC [21] 93.59 93.24 -0.35 52.9% 2.12
DMC [11] 93.62 92.69 -0.93 50.0% 2.00
GNN-RL [70] 93.49 93.59 0.10 54.0% 2.17
DepGraph w/o SL [9] 93.53 93.46 -0.07 52.6% 2.11
DepGraph with SL [9] 93.53 93.77 0.24 52.6% 2.11
ATO [62] 93.50 93.74 0.24 55.0% 2.22
Meta-Pruning (ours) 93.51 94.08 0.57 56.5% 2.30
C-SGD [7] 93.39 93.44 0.05 60.8% 2.55
Greg-1 [60] 93.36 93.18 -0.18 60.8% 2.55
Greg-2 [60] 93.36 93.36 0.00 60.8% 2.55
DepGraph w/o SL [9] 93.53 93.36 -0.17 60.2% 2.51
DepGraph with SL [9] 93.53 93.64 0.11 61.1% 2.57
ATO [62] 93.50 93.48 -0.02 65.3% 2.88
Meta-pruning (ours) 93.51 93.69 0.18 65.6% 2.91
Meta-pruning (ours) 93.51 93.25 -0.26 66.9% 3.02
Table 1 shows our results of pruning ResNet56 [16] on CIFAR10 [31]. Our algorithm outperforms all previous works. We can prune 65.6% FLOPs (i.e. 2.91x Speed Up) without any loss of accuracy, which is quite amazing. See Appendix B for more results and implementation details.
4.3 VGG19 on CIFAR100
Table 2: VGG19 on CIFAR100 (see full table in appendix 7)
Method Base Pruned $\Delta$ Acc Pruned FLOPs Speed Up
OBD [59] 73.34 60.66 -12.68 83.58% 6.09
EigenD [59] 73.34 65.18 -8.16 88.64% 8.80
Greg-2 [60] 74.02 67.75 -6.27 88.69% 8.84
DepGraph w/o SL [9] 73.50 67.60 -5.44 88.73% 8.87
DepGraph with SL [9] 73.50 70.39 -3.11 88.79% 8.92
Meta-Pruning (ours) 73.65 70.06 -3.59 88.85% 8.97
Table 2 shows our results of pruning VGG19 [54] on CIFAR100 [31]. Our algorithm reaches 70.06% pruned accuracy with 88.85% FLOPs pruned (i.e. 8.97x Speed Up). While our pruning speed up is larger than all previous works, our pruned accuracy is only a little bit lower than DepGraph with SL and higher than all other works. See Appendix C for more results and implementation details.
4.4 ResNet50 on ImageNet
Table 3: ResNet50 on ImageNet (see full table in appendix 10)
Method Base Top-1(Top-5) Pruned Top-1($\Delta$) Pruned Top-5($\Delta$) Pruned FLOPs
DCP [77] 76.01%(92.93%) 74.95%(-1.06%) 92.32%(-0.61%) 55.6%
CCP [49] 76.15%(92.87%) 75.21%(-0.94%) 92.42%(-0.45%) 54.1%
FPGM [19] 76.15%(92.87%) 74.83%(-1.32%) 92.32%(-0.55%) 53.5%
ABCP [42] 76.01%(92.96%) 73.86%(-2.15%) 91.69%(-1.27%) 54.3%
DMC [11] 76.15%(92.87%) 75.35%(-0.80%) 92.49%(-0.38%) 55.0%
Random [38] 76.15%(92.87%) 75.13%(-1.02%) 92.52%(-0.35%) 51.0%
DepGraph [9] 76.15%(-) 75.83%(-0.32%) - 51.7%
ATO [62] 76.13%(92.86%) 76.59%(+0.46%) 93.24%(+0.38%) 55.2%
DTP [39] 76.13%(-) 75.55%(-0.58%) - 56.7%
ours 76.14%(93.11%) 76.17%(+0.03%) 92.94%(-0.17%) 56.9%
Table 3 shows our results of pruning ResNet50 [16] on ImageNet[4]. Our algorithm reaches 76.17% top-1 and 92,95% top-2 accuracy with 56.9% FLOPs pruned. With the largest pruned FLOPs, our pruned accuracy is higher than almost all previous works (including DTP [39]). The only exception is that our accuracy is a bit lower than ATO [62], but our pruned FLOPs is much larger (1.7%) than it. See appendix D for more results and implementation details.
4.5 Ablation Study
How metanetwork works. To help readers visually grasp the effect of metanetwork, we visualized the "Test Acc VS. Speed Up" curve of both origin network and network after feedforward through metanetwork and finetuning . We generate the curve by prune the network little by little. After each pruning step, we evaluated the model’s performance on the test set (no finetuning in this step) and recorded a data point representing the current speed up and corresponding test accuracy. Connecting these points forms a curve that illustrates how accuracy changes as the speed up increases. The "Test Acc VS. Speed Up" curve of ResNet on CIFAR10 can be seen in figure 3(a). Similar curves of VGG19 on CIFAR100 and ResNet50 on ImageNet can be seen in appendix E.1.
As shown in the "Test Accuracy vs. Speed Up" curves, as the pruning speed up increases, the accuracy drops at a significantly slower rate after applying the metanetwork and finetuning. Take figure 3(a) as an example (this is exactly the metanetwork we use in 4.2). For the original network, the accuracy decreases rapidly and nearly collapses before reaching a speed up of 1.50x, suggesting that directly using group l2 norm criterion 3.3 to prune the origin network removes many important parts of the network (i.e. the origin network is hard to prune under this criterion). However, after metanetwork and finetuning, pruning to 2.0x speed up without finetuning causes almost no drop in accuracy, suggesting that using group l2 norm criterion on this network can efficiently distinguish the unimportant parts of this network (i.e. the network after metanetwork and finetuning is much easier to prune under this criterion). This supports our claim in 3 that metanetwork can transform a neural network that is difficult to prune into one that is much easier to prune.
(a) Origin VS After Metanetwork
(b) Impact Of Training Epochs On Metanetworks
Figure 3: Visualization of how metanetworks effect the networks (ResNet56 on CIFAR10)
Trade-off between accuracy and speed up. We observe an interesting phenomenon during meta-training: as the number of training epochs increases, the metanetwork’s ability to make the target network easier to prune becomes stronger. However, this comes at the cost of its ability to maintain the accuracy. As shown in figure 3(b), with more training epochs, the "Test Accuracy vs. Speed Up" curve shifts downward, and the flat portion of the curve becomes longer. This characteristic allows us to adaptively meet different pruning requirements. If a higher level of pruning is desired and a moderate drop in accuracy is acceptable, a metanetwork trained for more epochs would be preferable. Conversely, if preserving accuracy is more important, a metanetwork trained for fewer epochs would be a better choice.
Finetuning after Metanetwork. See appendix E.2 for our study on the impact of the number of finetuning epochs after metanetwork on performance.
5 Conclusion
In this paper we proposes a brand new meta-learning framework for network pruning. For the first time we introduce the concept of metanetwork (a network that takes another network as input and produces a modified network as output) into pruning. The metanetwork is trained to automatically learn effective pruning strategies. After training, it can take a network that is hard to prune as input and generate a transformed network that is much easier to prune. During pruning, our method needs nothing more than a feedforward through metanetwork and trivial finetuning to prune at state-of-the-art.
Advantages Over Previous Works. Our metanetwork is capable of considering the weights of multiple nearby layers, which sets it apart from most previous approaches. Unlike earlier methods that either focus only on local weights or incorporate more layers but rely on simple, hand-crafted relationships between them, our metanetwork automatically learns complex relationships among neighboring layers. These relationships are too intricate to be expressed by simple human-defined rules. So we use metanetwork to learn it automatically during training. This can explain why our work outperforms almost all previous works,
Limitations And Future Work. Our work also has several limitations. First, converting networks into graph representations requires GPUs with high VRAM, making it challenging to apply our method to large and densely parameterized models. Second, our current approach can’t be directly applied to networks with different architectures, such as Transformers or RNNs without appropriate modifications. However, these limitations also present opportunities for future research. We think the following directions are promising and worth exploring: (1) Extending our method to accommodate diverse architectures, including Transformers, diffusion models, and others. (2) Investigating alternative approaches to constructing the metanetworks. (3) Applying similar meta-learning frameworks to a broader range of tasks.
Acknowledgments and Disclosure of Funding
Use unnumbered first level headings for the acknowledgments. All acknowledgments go at the end of the paper before the list of references. Moreover, you are required to declare funding (financial activities supporting the submitted work) and competing interests (related financial activities outside the submitted work). More information about this disclosure can be found at: https://neurips.cc/Conferences/2025/PaperInformation/FundingDisclosure.
Do not include this section in the anonymized submission, only in the final paper. You can use the ack environment provided in the style file to automatically hide this section in the anonymized submission.
Appendix A General Experiment Setup
A.1 General Settings
Optimizer: For all data model training and fine-tuning, we use torch.optim.SGD with momentum=0.9. For meta-training, we employ torch.optim.AdamW. All learning rates are controlled using torch.optim.lr_scheduler.MultiStepLR with gamma=0.1. The learning rate is adjusted at specified milestones; for instance, if milestones=[50, 90], the learning rate is multiplied by 0.1 at epochs 50 and 90. These settings are not strictly necessary—other optimizers and schedulers may be equally effective.
Data for Meta-Training: As discussed in Section 3.4, meta-training requires two types of data: data models and traditional training datasets. To generate a data model, we train a base model, perform initial pruning followed by finetuning, and finally save the model along with relevant information. Typically, generating 3–10 data models is sufficient, with 2–8 used for meta-training and the remainder reserved for meta-evaluation or testing.
Meta Evaluation (Meta Eval): Meta evaluation is a process used to assess the quality of the metanetwork during meta-training. We did not include it in our final experiments due to its high computational cost; however, we retain it as an optional tool for future research. At the end of each meta-training epoch, we evaluate the metanetwork by feeding unseen data models (not used in meta-training) through the network, finetuning, and then pruning to achieve the maximum speed-up while maintaining accuracy above a predefined threshold. A higher resulting speed-up indicates better metanetwork performance.
Visualizing Metanetworks: This is a key technique used throughout our experiments. When we refer to visualizing a metanetwork, we mean passing a model through the metanetwork, followed by finetuning, and then visualize the “Test Accuracy vs. Speed-Up” curve of the resulting model, as illustrated in Figure 3. This visualization provides insights into the metanetwork’s behavior and helps determine whether it is suitable for pruning. An ideal metanetwork should exhibit a long flat region in the curve where accuracy remains close to or above the target pruning accuracy. There are two ways to select a suitable metanetwork: (1) using meta-evaluation during meta-training, or (2) visualizing metanetworks post-training. In our final experiments, we exclusively use the second method, which is significantly more efficient. We typically do not visualize all metanetworks but instead search for the best one using a binary search strategy.
Big Batch Size and Small Batch Size: We use two different batch sizes in our experiments: a small batch size and a big batch size. The big batch size is used during meta-training to compute accuracy loss on new networks. Due to memory constraints, such large batches cannot be processed in a single forward and backward pass. Instead, we accumulate gradients over multiple smaller mini-batches before performing a single optimizer update—a common PyTorch practice. The big batch size is thus expressed as batchsize * iters, indicating that one optimizer.step() is performed after iters forward/backward passes, each using a mini-batch of size batchsize. For all other standard training and finetuning tasks, we use the small batch size.
A.2 General Meta Training Details
Equivalent Conversion between Network and Graph: During meta-training, we first convert the original network into an origin graph. Then, we feed this origin graph through the metanetwork to generate a new graph. Finally, we convert the new graph back into a new network. When converting the original network into the origin graph, all network parameters are mapped to either node or edge features in the graph. However, some graph features do not correspond to any network parameter. For these missing features, we use default values that effectively simulate the absence of those parameters. When converting the new graph back into the new network, we only map those trainable parameters from the new graph back to the new network. For untrainable parameters in the new network, we keep it the same as the old network. We provide more explanations for this conversion process in the following subsections: B, C, and D.
The Relative Importance of Sparsity Loss and Accuracy Loss: In our experiments, we introduce a hyperparameter called "pruner reg", which controls the relative importance of sparsity loss compared to accuracy loss. During backpropagation, the gradient from the sparsity loss is scaled(multiplied) by this "pruner reg" value, while the gradient from the accuracy loss remains unscaled.
Meta Training Milestone: As discussed in the Ablation Study 3, as the number of training epochs increases, the metanetwork’s ability to make the network easier to prune becomes stronger, while its ability to maintain the accuracy becomes weaker. During meta-training, we usually set a milestone for our learning rate scheduler. Before this milestone, we use a relatively large learning rate to quickly improve the metanetwork’s ability to make networks more prunable. After the milestone, we switch to a smaller learning rate to finetune the metanetwork and slow down the change of the two abilities, allowing us to select a well-balanced metanetwork. To determine a reasonable milestone value, we can first perform a short meta-training phase using the large learning rate, then visualize the resulting metanetworks to identify the point where the accuracy-maintaining ability of metanetwork is a bit stronger than we expect.
Choose the Appropriate Metanetwork: During meta-training, we save the metanetwork after each epoch. After training is complete, we search for the most suitable metanetwork by visualizing its performance using a binary search strategy. Specifically, we start by visualizing the metanetwork from the middle epoch. If the accuracy of its flat region is below our target pruned accuracy, we next visualize the metanetwork from the first quarter of training. Otherwise, we check the one from the third quarter. We continue this process iteratively until we find a metanetwork that meets our criteria: a relatively high accuracy in the flat region and a long flat region indicating robustness to pruning.
A.3 General Pruning Details :
Pruning Speed Up: Once a suitable metanetwork has been selected for pruning, the next step is to determine the target speed up. In general, after visualizing the metanetwork, we observe a flat region followed by a sharp decline in accuracy. We define a turning point as the speed up at which the accuracy drops below a certain threshold. Experimentally, we find that we can safely prune the model at a speed up slightly higher than this turning point without incurring any accuracy loss after fine-tuning. For instance, when pruning ResNet-56 on CIFAR-10, we consider an accuracy drop below 0.93 as the turning point (typically around 2.0$\times$ speed up), and in practice, we are able to achieve a 2.9$\times$ speed-up with no loss in accuracy after finetuning.
Appendix B ResNet56 on CIFAR10
B.1 Equivalent Conversion between Network and Graph
We change ResNet56 into a graph with node featrues of 8 dimensions and edge features of 9 dimensions.
The node features consist of 4 features derived from the batch normalization parameters (weight, bias, running mean, and running variance) of the previous layer, along with 4 features from the batch normalization of the previous residual connection. If no such residual connection exists, we assign default values $[1,0,0,1]$ to the corresponding 4 features.
The edge features are constructed by zero-padding all convolutional kernels to a size of $3\times 3$ (note that our networks only contain kernels of size $3\times 3$ or $1\times 1$), and then flattening them into feature vectors. We treat all residual connections as convolutional layers with kernel size $1\times 1$. For example, consider two layers A and B connected via a residual connection, with neurons indexed as $1,2,3,\ldots$ in both layers. If the residual connection has no learnable parameters (i.e., it directly adds the input to the output), we represent the edge from $A_{i}$ to $B_{i}$ as a $1\times 1$ convolutional kernel with value 1, and the edge from $A_{i}$ to $B_{j}$ (where $i\neq j$) as a $1\times 1$ kernel with value 0. If the residual connection includes parameters (e.g., a downsample with a convolutional layer and batch normalization), we construct the edge features in the same way as for standard convolutional layers.
B.2 Meta Training
Data Models : We generate 10 models as our data models, among them, 8 are used for meta-training and 2 are used for validation (visualize the metanetwork). When generating each data model, we train them 200 epochs with learning rate 0.1 and milestone "100, 150, 180", then pruning with speed up 1.32x, followed by a finetuning for 80 epochs with learning rate 0.01 and milestone "40, 70". Finally, we can get a network with accuracy around 93.5%.
One Meta Training Epoch : Every epoch we enumerate over all 8 data models. When enumerate a data model, we feedforward it through the metanetwork and generate a new network. We feed all our CIFAR10 training data into the new network to calculate the accuracy loss, and use the parameters of new network to calculate the sparsity loss. Then we backward the gradients from both two losses to update our metanetwork.
Training: We train our metanetwork with learning rate 0.001, milestone "12", weight decay 0.0005 and pruner reg 10. Finally we use metanetwork from epoch 28 as our final network for pruning.
B.3 Full Results (Including Error Bars)
All results are summarized in Table 4, where we repeat the pruning process 3 times using different seeds: 7, 8, 9. It is important to note that due to the design of our algorithm and implementation, we cannot prune the network to exactly the same speed up across different runs. For instance, when targeting a 1.3$\times$ speed up, the actual achieved speed up may be slightly larger, such as 1.31$\times$, 1.32$\times$, or 1.35$\times$.
The aggregated statistical results are presented in Table 5, where each value is reported in the form of mean(standard deviation).
Table 4: ResNet56 on CIFAR10 Full
Method Base Pruned $\Delta$ Acc Pruned FLOPs Speed Up
NISP [69] 43.2% 1.76
Geometric [20] 93.59 93.26 -0.33 41.2% 1.70
Polar [76] 93.80 93.83 0.03 46.8% 1.88
DCP-Adapt [77] 93.80 93.81 0.01 47.0% 1.89
CP [35] 92.80 91.80 -1.00 50.0% 2.00
AMC [23] 92.80 91.90 -0.90 50.0% 2.00
HRank [41] 93.26 92.17 -1.09 50.0% 2.00
SFP [18] 93.59 93.36 -0.23 52.6% 2.11
ResRep [8] 93.71 93.71 0.00 52.8% 2.12
SCP [28] 93.69 93.23 -0.46 51.5% 2.06
FPGM [19] 93.59 92.93 -0.66 52.6% 2.11
FPC [21] 93.59 93.24 -0.35 52.9% 2.12
DMC [11] 93.62 92.69 -0.93 50.0% 2.00
GNN-RL [70] 93.49 93.59 0.10 54.0% 2.17
DepGraph w/o SL [9] 93.53 93.46 -0.07 52.6% 2.11
DepGraph with SL [9] 93.53 93.77 0.24 52.6% 2.11
ATO [62] 93.50 93.74 0.24 55.0% 2.22
Meta-Pruning (ours) 93.51 94.08 0.57 56.5% 2.30
Meta-Pruning (ours) 93.51 93.56 0.05 56.5% 2.30
Meta-Pruning (ours) 93.51 93.81 0.30 56.9% 2.32
GBN [67] 93.10 92.77 -0.33 60.2% 2.51
AFP [6] 93.93 92.94 -0.99 60.9% 2.56
C-SGD [7] 93.39 93.44 0.05 60.8% 2.55
Greg-1 [60] 93.36 93.18 -0.18 60.8% 2.55
Greg-2 [60] 93.36 93.36 0.00 60.8% 2.55
DepGraph w/o SL [9] 93.53 93.36 -0.17 60.2% 2.51
DepGraph with SL [9] 93.53 93.64 0.11 61.1% 2.57
ATO [62] 93.50 93.48 -0.02 65.3% 2.88
Meta-pruning (ours) 93.51 93.69 0.18 65.6% 2.91
Meta-pruning (ours) 93.51 93.42 -0.09 66.0% 2.94
Meta-pruning (ours) 93.51 93.31 -0.20 66.6% 2.99
Meta-pruning (ours) 93.51 93.25 -0.26 66.9% 3.02
Meta-pruning (ours) 93.51 93.20 -0.31 66.7% 3.00
Meta-pruning (ours) 93.51 93.18 -0.33 67.5% 3.08
Table 5: ResNet56 on CIFAR10 Statistics
Pruned Accuracy Pruned FLOPs Speed Up
93.82($\pm$0.21) 56.63%($\pm$0.19%) 2.3067($\pm$0.0094)
93.47($\pm$0.16) 66.07%($\pm$0.41%) 2.9467($\pm$0.0330)
93.21($\pm$0.03) 67.03%($\pm$0.34%) 3.0333($\pm$0.0340)
B.4 Computer Resources and Hyperparameters
See Table 6.
Table 6: ResNet56 on CIFAR10 Hyperparameters
Types Name Value
Compute Resources GPU NVIDIA RTX 4090
parallel No
Batch size small batch size 128
big batch size 500 $\times$ 100
Prepare Data Models data model num 10 ( 8 + 2 )
Train From Scratch epoch 200
lr 0.1
weight decay 0.0005
milestone "100, 150, 180"
Initial Pruning speed up 1.32
Finetuning epoch 80
lr 0.01
weight decay 0.0005
milestone "40, 70"
Metanetwork num layer 8
hiddim 64
in node dim 8
in edge dim 9
node res ratio 0.01
edge res ratio 0.01
Meta Training lr 0.001
weight decay 0.0005
milestone "12"
pruner reg 10
Final Pruning metanetwork epoch 28
speed up 2.3, 2.9, 3.0
Finetuning After Metanetwork epoch 100
lr 0.01
weight decay 0.0005
milestone "60, 90"
Finetuning After Pruning epoch 140
lr 0.01
weight decay 0.0005
milestone "80, 120"
Appendix C VGG19 on CIFAR100
C.1 Equivalent Conversion between Network and Graph
We change VGG19 into a graph with node featrues of 5 dimensions and edge features of 9 dimensions.
The node features consist of 4 features derived from the batch normalization parameters (weight, bias, running mean, and running variance) of the previous layer and 1 feature derived from the bias of previous layer (0 if bias doesn’t exist).
The edge features are constructed by zero-padding all convolutional kernels to a size of $3\times 3$ (note that our networks only contain kernels of size $3\times 3$ or $1\times 1$), and then flattening them into feature vectors.
C.2 Meta Training
Data Models : We generate 10 models as our data models, among them, 8 are used for meta-training and 2 are used for validation (visualize the metanetwork). When generating each data model, we train them 200 epochs with learning rate 0.1 and milestone "100, 150, 180", then pruning with speed up 2.0x, followed by a finetuning for 140 epochs with learning rate 0.01 and milestone "80, 120". Finally, we can get a network with accuracy around 73.5%.
One Meta Training Epoch : Every epoch we enumerate over all 8 data models. When enumerate a data model, we feedforward it through the metanetwork and generate a new network. We feed all our CIFAR100 training data into the new network to calculate the accuracy loss, and use the parameters of new network to calculate the sparsity loss. Then we backward the gradients from both two losses to update our metanetwork.
Training: We train our metanetwork with learning rate 0.001, milestone "15", weight decay 0.0005 and pruner reg 10. Finally we use metanetwork from epoch 51 as our final network for pruning.
C.3 Full Results (Including Error Bars)
All results are summarized in Table 7, where we repeat the pruning process 3 times using different seeds: 7, 8, 9. It is important to note that due to the design of our algorithm and implementation, we cannot prune the network to exactly the same speed up across different runs. For instance, when targeting a 8.95$\times$ speed up, the actual achieved speed up may be slightly larger, such as 8.95$\times$, 9.01$\times$, or 9.02$\times$.
The aggregated statistical results are presented in Table 8, where each value is reported in the form of mean(standard deviation).
Table 7: VGG19 on CIFAR100 Full
Method Base Pruned $\Delta$ Acc Pruned FLOPs Speed Up
OBD [59] 73.34 60.70 -12.64 82.55% 5.73
OBD [59] 73.34 60.66 -12.68 83.58% 6.09
EigenD [59] 73.34 65.18 -8.16 88.64% 8.80
Greg-1 [60] 74.02 67.55 -6.67 88.69% 8.84
Greg-2 [60] 74.02 67.75 -6.27 88.69% 8.84
DepGraph w/o SL [9] 73.50 67.60 -5.44 88.73% 8.87
DepGraph with SL [9] 73.50 70.39 -3.11 88.79% 8.92
Meta-Pruning (ours) 73.65 70.06 -3.59 88.85% 8.97
Meta-Pruning (ours) 73.65 69.19 -4.46 88.91% 9.02
Meta-Pruning (ours) 73.65 67.05 -6.60 89.07% 9.15
Table 8: VGG19 on CIFAR100 Statistics
Pruned Accuracy Pruned FLOPs Speed Up
68.77($\pm$1.26) 88.94%($\pm$0.09%) 9.0467($\pm$0.0759)
C.4 Computer Resources and Hyperparameters
See Table 9.
Table 9: VGG19 on CIFAR100 Hyperparameters
Types Name Value
Compute Resources GPU NVIDIA RTX 4090
parallel No
Batch size small batch size 128
big batch size 500 $\times$ 100
Prepare Data Models data model num 10 ( 8 + 2 )
Train From Scratch epoch 200
lr 0.1
weight decay 0.0005
milestone "100, 150, 180"
Initial Pruning speed up 2.0
Finetuning epoch 140
lr 0.01
weight decay 0.0005
milestone "80, 120"
Metanetwork num layer 8
hiddim 32
in node dim 5
in edge dim 9
node res ratio 0.01
edge res ratio 0.01
Meta Training lr 0.001
weight decay 0.0005
milestone "15"
pruner reg 10
Final Pruning metanetwork epoch 51
speed up 8.95
Finetuning After Metanetwork epoch 2000
lr 0.01
weight decay 0.0005
milestone "1850, 1950"
Finetuning After Pruning epoch 2000
lr 0.01
weight decay 0.0005
milestone "1850, 1950"
Appendix D ResNet50 on ImageNet
D.1 Equivalent Conversion between Network and Graph
We change ResNet50 into a graph with node featrues of 8 dimensions and edge features of 1, 9 or 49 dimensions. We employ 3 linear layers to project edge features of these 3 different dimensions into the same hidden dimension, enabling their integration into our graph neural network.
The node features consist of 4 features derived from the batch normalization parameters (weight, bias, running mean, and running variance) of the previous layer, along with 4 features from the batch normalization of the previous residual connection. If no such residual connection exists, we assign default values $[1,0,0,1]$ to the corresponding 4 features.
The edge features are constructed by simply flatten convolutional kernels into feature vectors (We have convolutional kernels of size $1\times 1$, $3\times 3$ or $7\times 7$). We treat all residual connections as convolutional layers with kernel size $1\times 1$. For example, consider two layers A and B connected via a residual connection, with neurons indexed as $1,2,3,\ldots$ in both layers. If the residual connection has no learnable parameters (i.e., it directly adds the input to the output), we represent the edge from $A_{i}$ to $B_{i}$ as a $1\times 1$ convolutional kernel with value 1. Different from section B, to reduce the VRAM footprint of the program, we build no edges between $A_{i}$ and $B_{j}$ ($i\neq j$). If the residual connection includes parameters (e.g., a downsample with a convolutional layer and batch normalization), we construct the edge features in the same way as for standard convolutional layers.
D.2 Meta Training
Data Models : We generate 3 models as our data models, among them, 2 are used for meta-training and 1 are used for validation (visualize the metanetwork). When generating each data model, we finetune them based on the pretrained weights for 30 epochs with learning rate 0.01 and milestone "10", then pruning with speed up 1.2920x, followed by a finetuning for 60 epochs with learning rate 0.01 and milestone "30". Finally, we can get a network with accuracy around 76.1%.
One Meta Training Epoch : We train with torch.nn.parallel.DistributedDataParallel (pytorch data parallel) across 8 gpus. Unlike in Section B and Section C, at the beginning of each epoch, we evenly distribute different data models across the GPUs. For example, since we have two data models, we load one on four GPUs and the other on the remaining four GPUs. At each iteration, we forward all eight models (replicated across the 8 GPUs) through the metanetwork to generate eight new models. We then feedforward a large batch of ImageNet data, evenly distributed through these eight new models, to compute the accuracy loss. We also use the parameters of the eight new networks to compute the sparsity loss. Then we backward the gradients from both two losses to update our metanetwork. Each training epoch consists of (ImageNet data num / big batch size) iterations, meaning that one full pass over the entire ImageNet dataset when computing the accuracy loss is considered as one epoch .
Training: We train our metanetwork with learning rate 0.01, milestone "4", weight decay 0.0005 and pruner reg 10. Finally we use metanetwork from epoch 13 as our final network for pruning.
D.3 Full Results (Including Error Bars)
All results are summarized in Table 10, where we repeat the pruning process 3 times. It is important to note that due to the design of our algorithm and implementation, we cannot prune the network to exactly the same speed up across different runs. For instance, when targeting a 2.3$\times$ speed up, the actual achieved speed up may be slightly larger, such as 2.31$\times$, 2.32$\times$, or 2.35$\times$.
The aggregated statistical results are presented in Table 11, where each value is reported in the form of mean(standard deviation).
Table 10: ResNet50 on ImageNet Full
Method Base Top-1(Top-5) Pruned Top-1($\Delta$) Pruned Top-5($\Delta$) Pruned FLOPs
DCP [77] 76.01%(92.93%) 74.95%(-1.06%) 92.32%(-0.61%) 55.6%
CCP [49] 76.15%(92.87%) 75.21%(-0.94%) 92.42%(-0.45%) 54.1%
FPGM [19] 76.15%(92.87%) 74.83%(-1.32%) 92.32%(-0.55%) 53.5%
ABCP [42] 76.01%(92.96%) 73.86%(-2.15%) 91.69%(-1.27%) 54.3%
DMC [11] 76.15%(92.87%) 75.35%(-0.80%) 92.49%(-0.38%) 55.0%
Random [38] 76.15%(92.87%) 75.13%(-1.02%) 92.52%(-0.35%) 51.0%
DepGraph [9] 76.15%(-) 75.83%(-0.32%) - 51.7%
ATO [62] 76.13%(92.86%) 76.59%(+0.46%) 93.24%(+0.38%) 55.2%
DTP [39] 76.13%(-) 75.55%(-0.58%) - 56.7%
ours 76.14%(93.11%) 76.17%(+0.03%) 92.94%(-0.17%) 56.9%
ours 76.14%(93.11%) 76.27%(+0.13%) 93.05%(-0.06%) 56.9%
ours 76.14%(93.11%) 76.15%(+0.01%) 92.95%(-0.16%) 56.9%
Table 11: ResNet50 on ImageNet Statistics
Pruned FLOPs (Speed Up) Pruned Top-1 Pruned Top-5
56.9%(2.32$\times$) 76.20%($\pm$0.05%) 92.98%($\pm$0.08%)
D.4 Computer Resources and Hyperparameters
See Table 12.
Table 12: ResNet50 on ImageNet Hyperparameters
Types Name Value
Compute Resources GPU NVIDIA A100 40G
parallel 8
Batch size small batch size 32 $\times$ 8
big batch size 32 $\times$ 8 $\times$ 200
Prepare Data Models data model num 3 ( 2 + 1 )
Generate Data Model epoch 30
lr 0.01
weight decay 0.0001
milestone "10"
Initial Pruning speed up 1.2920
Finetuning epoch 60
lr 0.01
weight decay 0.0001
milestone "30"
Metanetwork num layer 4
hiddim 11
in node dim 8
node res ratio 0.001
edge res ratio 0.001
Meta Training lr 0.01
weight decay 0.0005
milestone "4"
pruner reg 10
Final Pruning metanetwork epoch 13
speed up 2.3095
Finetuning After Metanetwork epoch 200
lr 0.01
weight decay 0.0001
milestone "120, 160, 185"
Finetuning After Pruning epoch 200
lr 0.01
weight decay 0.0001
milestone "120, 160, 185"
Appendix E More About Ablation Study
E.1 Visualize "Test Acc VS. Speed Up" Curves
As mentioned in section 4.5, we visualize the "Test Acc VS. Speed Up" curves of metanetworks in Resnet56 on CIFAR10 in figure 4, VGG19 on CIFAR100 in figure 5 and ResNet 50 on ImageNet in figure 6.
Figure 4: Resnet56 on CIFAR10 Visualization
Figure 5: VGG19 on CIFAR100 Visualization
Figure 6: ResNet50 on ImageNet Visualization
E.2 Finetuning after Metanetwork
(a) Finetune Different Epochs with Same Metanetwork
(b) Use Finetune to Get Better Performance
Figure 7: Finetune After Metanetwork (ResNet56 on CIFAR10)
As mentioned in section 3.5, our pruning pipeline can be summarized as:
$InitialPruning(Finetuning)\rightarrow Metanetwork(Finetuning)\rightarrow Pruning(Finetuning)$
In this pipeline, finetuning after pruning is a common way to improve the accuracy. However, finetuning after the metanetwork is first proposed by us and may raise questions for readers regarding its impact on performance—whether it is necessary or beneficial ? In this section, we aim to clarify the effects of finetuning after the metanetwork and demonstrate that it is both necessary and advantageous.
To illustrate this, we use the example of pruning ResNet56 on CIFAR10. In Figure 7(a), we visualize the "Test Accuracy vs. Speed Up" curves for models after the same metanetwork and after various amounts of finetuning. Without any finetuning after the metanetwork, the network’s accuracy is nearly zero (which is why there is no "finetune_0_epoch" line in the figure—it lies at the bottom). Directly proceeding to pruning and finetuning from such an unfinetuned network experimentally results in poor performance.
However, as we increase the number of finetuning epochs after the metanetwork, several interesting observations emerge:
  • Quickly Recover: The accuracy of our network after metanetwork quickly recovers after only very few finetuning, indicating that our metanetwork has the ability to preserve the accuracy (i.e. our accuracy loss in meta-training works)
  • Trend of Change: As the number of finetuning epochs increases, the flat portion of the curve becomes shorter, and the overall accuracy of this region improves.
When we feed the same network into metanetworks trained for different numbers of meta-training epochs and get several modified networks, then finetune them with different epochs to reach the same accuracy (as shown in Figure 7(b)), another important pattern emerges:
  • More Finetune Better Performance : A network obtained from a metanetwork with more meta-training epochs can typically achieve the same accuracy as one from a metanetwork with fewer meta-training epochs—but requires more finetuning. Importantly, the former usually results in better pruning performance. This is evident in Figure 7(b), where the green line (representing a higher meta-training epoch) exhibits a longer flat region, indicating improved robustness to pruning.
This suggests that the effectiveness of our method can be further improved by employing metanetworks trained for more meta-training epochs and applying additional finetuning afterward. In practice, however, we avoid excessive finetuning in order to keep the number of finetuning epochs within a reasonable range, allowing for multiple experimental trials.
Overall, all finetuning in our pipeline is necessary and effective.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The Abstract and Introduction sections accurately reflect the contribution and scope of the paper.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The limitations of our work is discussed in the conclusion section (5).
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: We propose no theoretical result that needs to be proved.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We open sourced our code and provide everything you need to reproduce our code in Appendix.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We provides a URL link to our released code. All information about datasets and how to run our code can be found together with our code.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: In Ablation Study and Appendix we talk about all the details of our experiments and how to choose hyparameter.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We run each experiment 3 times and calculate the error bars, which you can see in Appendix.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We report our computer resources for each experiment in Appendix
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: This paper conforms to the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: There is no societal impact of the work.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: This paper has no risks and needs no safeguards.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: The assets utilized in this paper are appropriately acknowledged. The associated license and terms of use have been clearly stated and duly adhered to.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We open sourced our code and model with clear explanations.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: We involve no crowdsourcing subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: We involve no crowdsourcing subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: We didn’t involve LLMs as any important, original, or non-standard components.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.